El vector G tiene un componente sobre el eje de X de 3 m, y un componente sobre el eje de Y de 5m.

The Vector G has a component above X axis of 3m, and a component above Y axis of 5m.

Answers:
a. 2.83 at an angle of 31 degrees
b. 14 at an angle of 59 degrees
c. 5.83 at an angle of 59 degrees
d. 34 at an angle of 72 degrees
Answer #1

Magnitude=sqrt(5^2+3^2)=5.83m

So magnitude=5.83 m

Tan^(-1)(5/3)=59degree

so option (c)is correct